The annual salaries of employees in a large company are approximately normally distributed with a mean of $50,000 and a standard deviation of $20,000.
a) What percent of people earn less than $40,000?
b) What percent of people earn between $45,000 and $65,000?
c) What percent of people earn more than $70,000?
1
Expert's answer
2020-11-18T19:40:15-0500

a.

"\mu=50,000"

"\sigma=20000"

"P(x<40000)=P(z<\frac{40000-50000}{20000})"

"=P(z<-0.5)=0.3085" from z tables.

Thus, 30.85% of people earn below $40,000

b.

P

P(45000<x<65000)

"P(45000<x<65000)=P(\frac{45000-50000}{20000}<z<\frac{65000-50000}{20000})"

"=P(-0.25<z<0.75)"

"=P(z<0.75)-P(z<-0.25)"

"=0.7734-0.4013" from z tables

"=0.3721"

37.21% of people earn between $45000 and $65000

c.

P(x>70000)

"P(x>70000)=1-P(z<\frac{70000-50000}{20000}"

"=1-P(z<1)"

=1-0.8413 from z tables

=0.1587

15.87% of the people earn more than $70,000.